# CITRAS: Covariate-Informed Transformer for Time Series Forecasting

### v1

@@ 4.1 Covariate-Informed Forecasting
EPF
NP
@@
0.079
0.206
0.071
0.199
0.071 0.197
0.133
0.259
0.092
0.225
0.108
0.233
0.101
0.233
0.193
0.327
4.2 Multivariate forecasting
For a comprehensive comparison, we further evaluate our model on well-recognized long-term multivariate forecasting datasets. To reflect practical scenarios, we adopt the rolling forecasting approach [23], where a model trained with predetermined input and forecasting lengths is evaluated with a longer forecasting length by iteratively integrating its output into the subsequent input. Following previous works [23, 24], we set the input length to 672, the forecasting length to {96, 192, 336, 720}, and the patch length to 96. We present the average results for the four forecasting lengths in Table 2. CITRAS ranks first or second in most benchmarks, highlighting the effectiveness and versatility of its design.
Table 2: Results of the multivariate forecasting, where the input length is set to 672, and the forecasting length is set to {96, 192, 336, 720}. We adopt the rolling forecasting approach following [23] where one model is used for four forecasting lengths. Averaged results are reported here and full results are provided in Table 5. Results are cited from Timer-XL [24] if available; otherwise reproduced.
Model
CITRAS
Timer-XL
TimeXer
iTrans.
TFT
Cross.
TimesNet
Auto.
Metric
@@ 4.2 Multivariate forecasting
ETTh1
@@
0.240
0.273
0.223
0.263
0.266
0.291
0.308
0.333
0.225
0.286
0.264
0.293
0.435
0.455
Traffic
0.379
0.254
0.374
0.255
0.401
0.288
0.384
0.274
0.590
0.322
0.513
0.294
0.602
0.322
0.684
0.433
4.3 Ablation Study
Table 3: Ablation Results of Attention Score Smoothing
Setting
w/o
w/
@@ 4.3 Ablation Study
0.074
0.070 (-4.3%)
ETTh1
0.401
0.389 (-3.2%)
ECL
0.165
0.160 (-3.2%)
CITRAS adopts two novel mechanisms: KV Shift and Attention Score Smoothing. To clarify the effectiveness of KV Shift, as well as the importance of covariates, we trained and tested our model under three covariate usage settings: None (no covariates used), P (observed and past known covariates used), and P+F (P and future known covariates used). Figure 3 demonstrates that forecasting solely relying on the target has limitations, and covariates play a crucial role in performance improvement. In particular, future known covariates provide much richer context extending into the forecasting horizon, highlighting the indispensability of KV Shift to incorporate them. Furthermore, compared to other baselines, CITRAS shows a higher rate of performance improvement through the use of covariates, supporting its suitability for covariate-informed forecasting.
We also trained and tested an ablated version of our model without Attention Score Smoothing. Table 3 shows that Attention Score Smoothing is effective in both covariate-informed and multivariate settings. This supports the versatility of this mechanism, which captures variate-level cross-variate dependencies by smoothing over detailed patch-level dependencies.
Figure 3: Comparison of MSE on the EPF-NP and EDF datasets under three covariate usage settings: None (no covariates used), P (observed and past known covariates used), and P+F (P and future known covariates used). The MSE change rates shown on the bar graph are calculated with respect to the None setting as the baseline.
5 Conclusion
Considering the critical roles of covariates in practical time series forecasting, we propose CITRAS, a patch-based Transformer that flexibly leverages multiple targets, observed covariates, and known covariates. Specifically, we introduce two novel mechanisms within the patch-wise cross-variate attention modules: Key-Value (KV) Shift, which seamlessly integrates future known covariates with other variables, and Attention Score Smoothing, which transforms locally accurate dependencies into global variate-level dependencies. Experimentally, CITRAS achieves state-of-the-art performance and demonstrates strong cross-variate understanding in both covariate-informed and multivariate settings. In the future, we will explore foundational temporal understanding capability in covariate-informed settings.
Appendix A Implementation Details
A.1 Dataset Descriptions
We perform short-term covariate-informed forecasting on the electricity price forecasting (EPF) datasets and the electricity demand forecasting (EDF) dataset. The EPF datasets [18] consist of five subsets from different day-ahead electricity markets, each spanning six years: (1) EPF-NP documents the Nord Pool electricity market from 2013-01-01 to 2018-12-24, containing hourly electricity prices as a target variable, with corresponding grid load and wind power forecasts as known covariates. (2) EPF-PJM records the Pennsylvania-New Jersey-Maryland market from 2013-01-01 to 2018-12-24, containing zonal electricity prices in the Commonwealth Edison (COMED) as a target variable, with corresponding system load and COMED load forecasts as known covariates. (3) EPF-BE captures Belgium’s electricity market from 2011-01-09 to 2016-12-31, containing hourly electricity prices as a target variable, with corresponding load forecasts in Belgium and generation forecasts in France as known covariates. (4) EPF-FR documents the electricity market in France from 2012-01-09 to 2017-12-31, containing hourly prices as a target variable, with corresponding load and generation forecasts as known covariates. (5) EPF-DE records the German electricity market from 2012-01-09 to 2017-12-31, containing hourly prices as a target variable, with zonal load forecasts in the TSO Amprion zone and wind and solar generation forecasts as known covariates. (6) The EDF dataset [2] is an hourly time series dataset spanning over five years from 2015-01-03 to 2020-06-27, where the target variable is the national electricity demand in Panama. Observed covariates include four meteorological indices (temperature, relative humidity, liquid precipitation, wind speed) measured in three cities (Tocumen, Santiago, David) in Panama, resulting in a total of 12 indices. In addition, we adopt three binary calendar indicators (holiday, weekend, school day) as known covariates.
In addition, we perform long-term multivariate forecasting on 7 real-world datasets: (1) ETT [44] contains 7 monitoring factors in the electricity transformers from July 2016 to July 2018. ETTh1 and ETTh2 are hourly subsets, and ETTm1 and ETTm2 are subsets recorded every 15 minutes. (2) Weather [40] includes 21 meteorological factors recorded every 10 minutes from the Weather Station of the Max Planck Biogeochemistry Institute in 2020. (3) ECL [40] comprises hourly electricity consumption data from 321 clients. (4) Traffic [40] records hourly road occupancy rates measured by 862 sensors on San Francisco Bay area freeways from January 2015 to December 2016. We use all the variables in these datasets as target variables. The details of datasets are provided in Table 4.
EPF-NP 1 0 2 1 Hour (36500, 5219, 10460)
EPF-PJM 1 0 2 1 Hour (36500, 5219, 10460)
Covariate-Informed EPF-BE 1 0 2 1 Hour (36500, 5219, 10460)
Forecasting EPF-FR 1 0 2 1 Hour (36500, 5219, 10460)
EPF-DE 1 0 2 1 Hour (36500, 5219, 10460)
EDF 1 12 3 1 Hour (33442, 4783, 9586)
ETTh 7 0 0 1 Hour (8545, 2881, 2881)
Multivariate ETTm 7 0 0 15 Minutes (34465, 11521, 11521)
Forecasting ECL 321 0 0 1 Hour (18317, 2633, 5261)
Weather 21 0 0 10 Minutes (36792, 5271, 10540)
Traffic 862 0 0 1 Hour (12185, 1757, 3509)
A.2 Experiment settings
All the experiments are implemented in PyTorch [30]. We utilize a single NVIDIA A100 80GB GPU for the ECL and Traffic datasets and a single NVIDIA V100 32GB GPU for the other datasets. We adopt ADAM [17] with an initial learning rate of $10^{-4}$ and L2 loss for model optimization. The training process is fixed at 10 epochs with early stopping. We set the number of layers in our proposed model $L\in\{1,2,3,4\}$, the number of heads in the multi-head attention layers as 8, dimension of token embedding $D\in\{128,256,512,1024\}$, and a smoothing factor for Attention Score Smoothing $\alpha\in\{0.1,0.2\}$. The patch length is uniformly set to 24 for short-term covariate-informed forecasting and 96 for long-term multivariate forecasting.
A.3 Baseline Descriptions
We cite the baseline results in Table 1 from the TimeXer [34] paper and the results in Tables 2 and 5 from the Timer-XL [24] paper, if available. Otherwise, we reproduced the baseline models based on the TimesNet [39] repository. Note that Timer-XL, TimeXer, and iTransformer are extended from their original implementations to support future known covariates in short-term covariate-informed forecasting. Specifically, for Timer-XL, we modified the masking scheme in TimeAttention so that target token embeddings can attend to known covariate token embeddings from one patch step ahead. For TimeXer and iTransformer, we introduced an additional variate embedding layer for known covariates to handle their longer lengths. On the other hand, Crossformer, TimesNet, and Autoformer remain in their original implementations, which cannot accommodate future covariates as the extension is not straightforward.
Appendix B Full Results
B.1 Full Results of Multivariate forecasting

### v2

@@ 4.2 Multivariate Forecasting
ETTh1
0.389
0.421
0.428
0.453
0.409
0.430
0.421
0.445
0.406
0.431
0.414
0.433
0.413
0.432
0.415
0.426
0.426
0.444
ETTh2
0.359
0.405
0.353
0.405
0.372
0.409
0.383
0.415
0.339
0.391
0.340
0.388
0.336
0.388 0.339
0.386
0.352
0.404
ETTm1
0.350
0.382
0.368
0.394
0.368
0.390
0.377
0.404
0.358
0.384
0.355
0.381
0.369
0.393
0.358
0.379
0.357
0.380
ETTm2
0.262
0.320
0.266
0.321
0.275
0.327
0.278
0.334
0.254
0.315
0.247
0.308
0.267
0.326
0.254 0.314
0.262
0.326
ECL
0.160 0.253
0.166
0.264
0.155
0.246
0.164
0.258
0.161
0.253 0.160
0.255
0.183
0.291
0.171
0.269
0.165
0.265
Weather
0.233
0.272
0.223
0.263
0.240
0.273
0.266
0.291
0.225
0.265
0.221
0.266
0.232
0.273
0.229
0.265
0.239
0.291
Traffic
0.379
0.254
0.401
0.288
0.374
0.255
0.384
0.274
0.386
0.270
0.400
0.280
0.446
0.327
0.415
0.286
0.423
0.298
PEMS04
0.126
0.220
0.147
0.261
0.126
0.227
0.135
0.237
0.126
0.228
0.186
0.285
0.141
0.261
0.262
0.340
0.244
0.333
PEMS08
0.236
0.216
0.270
0.279
0.247 0.234
0.280
0.253
0.265
0.244
0.330
0.321
0.337
0.324
0.442
0.378
0.408
0.380
4.3 Model Analysis
4.3.1 Effect of Covariates
To illustrate the importance of covariates, we evaluate the performance of three top-ranking models, CITRAS, Timer-XL, and TimeXer, under three covariate usage settings: None (no covariate used), P (past covariates used), and P+F (past and future covariates used). Figure 3 demonstrates that relying solely on the target variable limits forecasting performance across all models, while covariates, particularly future covariates, play a crucial role in enhancing performance. Among these models, CITRAS exhibits the highest rate of improvement through the use of covariates, highlighting its effectiveness for covariate-informed forecasting. Visualization results that intuitively demonstrate the importance of covariates are in Appendix C.
Figure 3: Comparison of MSE under three variable usage settings: None (no covariate used), P (past covariates used), and P+F (past and future covariates used). The MSE change rates shown on the bar graph are calculated with respect to the "None" setting of each model.
4.3.2 Ablation Study
To elucidate the strong covariate understanding capability of CITRAS, we conduct the following ablation studies.
KV Shift
KV Shift captures the dependency between target variables (query) and known covariates (key) at the concurrent patch step $i$, and incorporates future known covariates (value) at step $i+1$. Another straightforward method to utilize future known covariates is to drop the first patch of known covariates and left-shift this channel entirely (denoted as Variate Shift), thereby evaluating the dependency between target variable at step $i$ (query) and future known covariates at step $i+1$ (key). However, Variate Shift imposes an additional burden on the model to reconcile the step misalignment in capturing the correlation, resulting in inferior performance compared to KV Shift, as shown in Table 3.
Table 3: Ablation Results of KV Shift. "Patch Step" represents the step of the key and value of the known covariate corresponding to the query of patch step $i$.
Design
Patch Step
EPF-NP
EDF
BS
Key
Value
MSE
MAE
MSE
MAE
MSE
MAE
w/o future
$i$
$i$
0.226
0.259
0.086
0.205
0.289
0.320
Variate Shift
$i+1$
$i+1$
0.176
0.218
0.071
0.194
0.284
0.319
KV Shift
$i$
$i+1$
@@ 4.3.2 Ablation Study
0.280
0.313
Attention Score Smoothing
Attention Score Smoothing captures locally accurate patch-level cross-variate dependency and refines it into global variate-level dependency. Note that a large alpha emphasizes patch-level dependency (with $\alpha=1.0$ corresponding to no smoothing), while a smaller alpha emphasizes variate-level dependency through heavier smoothing. Figure 4 (Left) illustrates the performance at different levels of granularity in cross-variate attention. The performance of the variate-level approach is represented by TimeXer. We observe that the suitability of variate-level versus patch-level approaches is dependent on the dataset. However, strong Attention Score Smoothing (around $\alpha=0.2$) strikes a balance between the two, highlighting the versatile effectiveness of this mechanism in both covariate-informed and multivariate forecasting.
4.3.3 Representation Analysis
To uncover the effects of Attention Score Smoothing further, we present the visualization of the cross-variate attention scores between the target variable and other variables in the EDF dataset in Figure 4 (Right). It is observed that patch-wise attention scores effectively capture the local negative impact of "Holiday" on the target variable ("Electricity Demand") . However, these scores are also susceptible to local disturbances, as demonstrated by the relatively fluctuating scores of "Humidity" compared to the "Temperature" that shows stable correlation with the target variable. By applying Attention Score Smoothing, the obtained scores effectively convey the strong dependency on "Holiday" from past steps and mitigate the noisy interaction caused by the fluctuating scores of "Humidity". In this way, Attention Score Smoothing enjoys the advantages of both patch-level and variate-level approaches while offering improved interpretability.
Figure 4: Left: The performance at different levels of granularity in cross-variate attention. The performance at "Variate" level is represented by TimeXer. Strong Attention Score Smoothing (around $\alpha=0.2$) outperforms both variate-level and patch-level approach. Right: Visualization of attention scores between the target variable ("Electricity Demand") and other variables observed in EDF dataset. By applying Attention Score Smoothing, the smoothened scores capture strong dependency on "Holiday" while mitigating noisy interaction between "Humidity".
5 Conclusion
Considering the critical roles of covariates in practical time series forecasting, we propose CITRAS, a decoder-only Transformer that flexibly leverages multiple targets, observed covariates, and known covariates. Specifically, we introduce two novel mechanisms within the patch-wise cross-variate attention modules: Key-Value (KV) Shift, which seamlessly integrates future values of known covariates into forecasting, and Attention Score Smoothing, which refines locally accurate dependencies into global variate-level dependencies. Experimentally, CITRAS outperforms state-of-the-art models on thirteen benchmarks from both covariate-informed and multivariate settings, demonstrating its versatile ability to leverage cross-variate and cross-time dependencies for improved forecasting accuracy. In the future, we will explore the generalization capability of CITRAS through large-scale pre-training to take advantage of its simple decoder-only Transformer architecture.
Appendix A Implementation Details
A.1 Dataset Descriptions
We perform covariate-informed forecasting on seven real world datasets. EPF datasets [23] consist of five subsets from different day-ahead electricity markets, each spanning six years: (1) EPF-NP documents the Nord Pool electricity market from 2013-01-01 to 2018-12-24, containing hourly electricity prices as a target variable, with corresponding grid load and wind power forecasts as known covariates. (2) EPF-PJM records the Pennsylvania-New Jersey-Maryland market from 2013-01-01 to 2018-12-24, containing zonal electricity prices in the Commonwealth Edison (COMED) as a target variable, with corresponding system load and COMED load forecasts as known covariates. (3) EPF-BE captures Belgium’s electricity market from 2011-01-09 to 2016-12-31, containing hourly electricity prices as a target variable, with corresponding load forecasts in Belgium and generation forecasts in France as known covariates. (4) EPF-FR documents the electricity market in France from 2012-01-09 to 2017-12-31, containing hourly prices as a target variable, with corresponding load and generation forecasts as known covariates. (5) EPF-DE records the German electricity market from 2012-01-09 to 2017-12-31, containing hourly prices as a target variable, with zonal load forecasts in the TSO Amprion zone and wind and solar generation forecasts as known covariates. (6) EDF dataset [2] is an hourly time series dataset spanning over five years from 2015-01-03 to 2020-06-27, where the target variable is the national electricity demand in Panama. Observed covariates include four meteorological indicators (T2M=Air temperature at 2 m (${}^{\circ}\mathrm{C}$), QV2M=Specific humidity at 2 m (%), TQL=Total precipitable liquid water ($1/m^{2}$), and W2M=Wind speeed at 2 m ($m/s$)) measured in three cities (Tocumen, Santiago, David) in Panama, resulting in a total of 12 indices. In addition, we adopt three binary calendar features (holiday, school day, weekend) as known covariates. (7) BS [17] is an hourly time series data spanning two years from 2011-01-01 to 2012-12-31, where the target variables are three types of rental counts (casual, registered, and cnt=total count) in Bike Sharing system in Washington D.C. Observed covariates include five meteorological indicators (weathersit=weather situation, temp=temperature, atemp=feeling temperature, hum=humidity, and windspeed). In addition, we adopt three calendar features (holiday, weekday, working day) as known covariates.
In addition, we perform multivariate forecasting on nine real-world datasets: ETT [56] contains 7 monitoring factors in the electricity transformers from July 2016 to July 2018. (1) ETTh1 and (2) ETTh2 are hourly subsets, and (3) ETTm1 and (4) ETTm2 are subsets recorded every 15 minutes. (5) Weather [49] includes 21 meteorological factors recorded every 10 minutes from the Weather Station of the Max Planck Biogeochemistry Institute in 2020. (6) ECL [49] comprises hourly electricity consumption data from 321 clients. (7) Traffic [49] records hourly road occupancy rates measured by 862 sensors on San Francisco Bay area freeways from January 2015 to December 2016. (8) PEMS04 [25] contains the public traffic network data recorded at 307 locations in California every 5 minutes from 2018-01-01 to 2018-02-28. Similarly, (9) PEMS08 [25] contains the data at 170 locations from 2016-07-01 to 2016-08-31. We use all the variables in these datasets as target variables. The summary of datasets are provided in Table 4.
EPF-NP 1 0 2 1 Hour (36500, 5219, 10460)
EPF-PJM 1 0 2 1 Hour (36500, 5219, 10460)
Covariate-Informed EPF-BE 1 0 2 1 Hour (36500, 5219, 10460)
Forecasting EPF-FR 1 0 2 1 Hour (36500, 5219, 10460)
EPF-DE 1 0 2 1 Hour (36500, 5219, 10460)
EDF 1 12 3 1 Hour (33442, 4783, 9586)
BS 3 5 3 1 Hour (11974, 1716, 3452)
ETTh 7 0 0 1 Hour (8545, 2881, 2881)
ETTm 7 0 0 15 Minutes (34465, 11521, 11521)
Multivariate ECL 321 0 0 1 Hour (18317, 2633, 5261)
Forecasting Weather 21 0 0 10 Minutes (36792, 5271, 10540)
Traffic 862 0 0 1 Hour (12185, 1757, 3509)
PEMS04 307 0 0 5 Minutes (10100, 3400, 3399)
PEMS08 170 0 0 5 Minutes (10618, 3573, 3572)
A.2 Experiment Settings
All the experiments are implemented in PyTorch [36]. We utilize a single NVIDIA A100 80GB GPU for the ECL, Traffic, and PEMS datasets and a single NVIDIA V100 32GB GPU for the other datasets. We apply Series Stationarization [31] to the embedding and projection layers of CITRAS to mitigate nonstationarity. We adopt ADAM [22] with an initial learning rate of $10^{-4}$ and L2 loss for model optimization. The training process is fixed at 10 epochs with an early stopping mechanism, which terminates training if the validation performance does not improve for 3 consecutive epochs. We set the number of layers in our proposed model $L\in\{1,2,3,4\}$, the number of heads in the multi-head attention layers as 8, dimension of token embedding $D\in\{128,256,512,1024\}$, and the smoothing factor for Attention Score Smoothing $\alpha\in\{0.1,0.2,0.4,0.8\}$. The patch length is uniformly set to 24 for covariate-informed forecasting and 96 for multivariate forecasting. We ensure that the drop last option is set to False in the test data loader to avoid any unfair impact, as highlighted in [37]. For the ablation study, we use the same hyperparameters as the full model to ensure a fair comparison.
A.3 Baseline Descriptions
We reproduced the baseline models based on the TimesNet [48] repository or official repositories. Note that Timer-XL, TimeXer  [^1] The TimeXer implementation that utilizes future covariates is not officially released., iTransformer, and Leddam are extended from their original implementations to support future covariates in covariate-informed forecasting. Specifically, for Timer-XL, we modified the masking scheme in TimeAttention so that target token embeddings can attend to known covariate token embeddings from one patch step ahead. For TimeXer, iTransformer, and Leddam, we introduced an additional variate embedding layer for known covariates to handle their longer lengths.
Appendix B Extra Model Analysis
B.1 Comparison between Decoder-only and Encoder-only Designs
CITRAS achieves the integration of covariates while preserving the canonical decoder-only Transformer architecture. This is advantageous because this design maintains the causality of forecasting and token-wise supervision, thereby realizing strong autoregressive capabilities, as pointed out in  [30]. To examine the benefits of the decoder-only approach, we designed an encoder-only variation of CITRAS. In this variation, we remove causal masking in cross-time attention, allowing each token to attend to others along the time dimension. Additionally, we apply a projection layer to the flattened features of all tokens in the target variable to predict future values. Table 5 and Table 6 present the comparison results between the encoder-only and decoder-only designs in covariate-informed settings (with future covariates) and multivariate settings, respectively. It is observed that in both settings, the decoder-only design outperforms the encoder-only design across most datasets. This result demonstrates the strength of CITRAS, which seamlessly integrates covariates into the decoder-only Transformer to effectively model both cross-variate and cross-time dependencies for improved forecasting accuracy.
Table 5: Comparison of encoder-only and decoder-only designs on the covariate-informed forecasting
Design
EPF-NP
EPF-PJM
EPF-BE
EPF-FR
EPF-DE
EDF
BS
@@ B.1 Comparison between Decoder-only and Encoder-only Designs
Encoder-only
0.188
0.236
0.084
0.181
0.365
0.240
0.399
0.198
0.240
0.312
0.068
0.189
0.366
0.376
Decoder-only (ours)
0.171
0.214
0.063
0.153
0.362
0.248
0.363
0.176
0.217
0.293
0.070
0.193
0.280
0.313
Table 6: Comparison of encoder-only and decoder-only designs on the multivariate forecasting. The evaluation protocol is the same as in Table 2, and averaged results across the four forecasting lengths {96, 192, 336, 720} are shown.
Design
ETTh1
ETTh2
ETTm1
ETTm2
ECL
Weather
PEMS08
@@
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Encoder-only
0.428
0.446
0.371
0.410
0.366
0.397
0.288
0.342
0.162
0.254
0.227
0.268
0.257
0.240
Decoder-only (ours)
@@
0.236
0.216
B.2 Learnable Smoothing Factor
The smoothing factor ($\alpha$) in Attention Score Smoothing can be treated as a learnable parameter. However, as shown in Table 7, we empirically find that a fixed smoothing factor is sufficient to achieve the desired smoothing effect. Consequently, our Attention Score Smoothing introduces no new learning parameters.
Table 7: Comparison between learnable and fixed smoothing factor
Smoothing factor
EPF-NP
EDF
ETTm1
ETTm2
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Learnable
0.174
0.218
0.070
0.193
0.356
0.383
0.267
0.324
Fixed (ours)
0.171
0.214
0.070
0.193
0.350
0.382
0.262
0.320
Appendix C Showcase
To visually demonstrate the covariate understanding capability of CITRAS, we present visualization results for the EPF-NP, EDF, and BS datasets in Figure 5. We show the forecasting results under two covariate usage settings: None (no covariates used) and P+F (past and future covariates used). It is observed that CITRAS can capture both positive and negative correlations between the target variable and covariates effectively. In the EPF-NP result, CITRAS accurately predicts the sudden rise in target electricity prices by considering the high grid load during the forecasting horizon. In the EDF result, it adjusts the electricity demand downward by recognizing that the forecasting horizon falls on a holiday. In the BS result, CITRAS effectively captures the negative correlation between working days and bike rental counts, accurately forecasting the rise in counts on the day following a working day. Interestingly, patch-based models, including CITRAS and Timer-XL, are better at reflecting the local impact of future covariates on target forecasting than TimeXer. This is likely because the variate-level approach adopted by TimeXer cannot accurately capture the local impact of covariates within the forecasting horizon, as it combines these pivotal features and past values into a single variate token.
Figure 5: Visualization results for the EPF-NP (top), EDF (middle), and BS (bottom) datasets under two covariate usage settings: None (no covariates used, colored in green) and P+F (past and future covariates used, colored in blue). Observed covariates are omitted from the visualization.
Appendix D Computational Efficiency
Consider the covariate-informed setting where the number of target variables is $1$, the number of covariates (including observed and known covariates) is $C$, the patch length of the target variable is $N$. CITRAS calculates cross-variate attention between the target and covariates at each patch step, resulting in a computational complexity of $O(CN)$. Additionally, it applies cross-time attention to each variate separately, incurring a complexity of $O(CN^{2})$. This is favorable compared to the Any-variate Attention adopted by Timer-XL, which incurs $O(C^{2}N^{2})$ to flexibly accommodate covariates. Furthermore, the $O(CN)$ complexity of cross-variate attention is advantageous when the number of covariates $C$ increases, compared to iTransformer, which increases to $O(C^{2})$. We acknowledge that CITRAS has quadratic complexity in the time dimension. Also, the complexity of cross-variate attention can be quadratic with respect to the number of target variables. However, we can introduce advanced attention modules, such as linear attention, to reduce the complexity, as it is based on the simple Transformer architecture.
Besides the theoretical analysis, we demonstrate the empirical model efficiency on the BS dataset and PEMS08 dataset. We use the best model configuration for each model and evaluate using the same batch size for a fair comparison. The results in the BS dataset underscore the efficient model design of CITRAS. In the PEMS08 dataset, we observe the trade-off between training time and model performance in terms of MSE. Among them, CITRAS achieves the best performance without significant compromise in efficiency.
Figure 6: Comparison of model efficiency on the BS and PEMS08 datasets. Note that the x-axis is in logarithmic scale for the PEMS08 result.
Appendix E Error Bar
We report the standard deviation of CITRAS performance on covariate-informed forecasting with future known covariates under five runs with different random seeds. We also refer to the result of some competitive baselines from Table 1. The result in Table 8 exhibits that the performance of CITRAS is stable.
Table 8: Error bar of CITRAS on the covariate-informed forecasting with future covariates. The results of baselines are from Table 1 for reference.
Model
CITRAS
TimeXer
Timer-XL
iTrans.
@@ Appendix E Error Bar
MAE
MSE
MAE
MSE
MAE
EPF–NP
0.171
$\pm$0.004
0.214
$\pm$0.004
0.175
0.225
0.176
0.220
0.203
0.250
EPF-PJM
0.0625
$\pm$0.0013
0.153
$\pm$0.001
0.0796
0.175
0.0623
0.155
0.0773
0.171
EPF–BE
0.362
$\pm$0.010
0.248
$\pm$0.004
0.368
0.231
0.384
0.257
0.339
0.246
EPF–FR
0.363
$\pm$0.010
0.176
$\pm$0.002
0.364
0.193
0.430
0.214
0.378
0.205
EPF–DE
0.217
$\pm$0.003
0.293
$\pm$0.001
0.283
0.322
0.228
0.299
0.258
0.327
EDF
0.0704
$\pm$0.0005
0.193
$\pm$0.001
0.0714
0.199
0.0791
0.206
0.0712
0.197
BS
0.280
$\pm$0.001
0.313
$\pm$0.001
0.338
0.357
0.286
0.318
0.358
0.360
Appendix F Extra Results
F.1 Comparison with Pre-trained Models
We further compare CITRAS with large pre-trained models under covariate-informed settings with future covariates. For baselines, we choose Moirai [46] (moirai-1.1-R-large) and Moirai-MoE [27] (moirai-moe-1.0-R-base), which are state-of-the-art pre-trained models that natively support future covariates. For a fair comparison, we present the zero-shot results of these models and the full-shot results of Moirai, which is fine-tuned with the full training dataset for a maximum of five epochs with a learning rate $lr\in\{1e-5,1e-6,1e-7\}$. Table 9 shows that CITRAS consistently outperforms Moirai under full-shot settings. Note that the implementation of ChronosX [4], which fine-tunes univariate pre-trained models to integrate covariates, is not publicly available yet and therefore not comparable.
Table 9: Comparison with pre-trained models on the covariate-informed forecasting
Setting Model
@@ F.1 Comparison with Pre-trained Models
BS
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
full
CITRAS
0.171
0.214
0.063
0.153
0.362
0.248
0.363
0.176
0.217
0.293
0.070
0.193
0.280
0.313
shot
Moirai
0.258
0.270
0.083
0.184
0.372
0.253
0.369
0.209
0.462
0.428
0.131
0.247
0.389
0.353
zero
Moira-MoE
0.260
0.279
0.076
0.172
0.362
0.251
0.370
0.204
0.500
0.451
0.196
0.316
0.684
0.496
shot
Moirai
0.294
0.296
0.086
0.186
0.393
0.271
0.398
0.220
0.544
0.476
0.144
0.268
0.554
0.413
F.2 Full Results of Multivariate Forecasting